write the general equation form of the line which passes through the 2 points: (1,2), (6,7)
Found 2 solutions by tutorcecilia, Edwin McCravy:
Answer by tutorcecilia(2144)

Ax+By=C [General equation of a line]
(y-y2)=m(x-x2) [Point-slope equation of a line.]
m=(y-y2)/(x-x2) [Formula for the slope of a line.]
.
.
(1,2), (6,7) [Find the slope of the line with these two points]
m=(7-2)/(6-1)=5/5=1 [Plug-in the values]
.
(y-7)=(1)(x-6) [Use the point-slope formula; plug-in the points and slope]
y-7=x-6 [Isolate the numbers]
y-x=-6+7
-x+y=1 [General equation of a line]
.
Check by plugging either set of coordinate points into the formula:
Example: Use (1, 2)
(-(1)+(2))=1
1=1 [checks out]
.
Use (6, 7)
-(6)+7=1
1=1 [Also checks out]

Answer by Edwin McCravy(6938)

write the general equation form of the line 
which passes through the 2 points: (1,2), (6,7)

You need two formulas:

     y2 - y1
m = --------- to find the slope, m, and
     x2 - x1

then you substitute the value you get for m, then 
substitute the values for x1 and y1. Do not
substitute anything for x and y.  Just leave them
variable:

y - y1 = m(x - x1)

(x1,y1) = (1,2), (x2, y2) = (6,7)

so

     y2 - y1     7 - 2     5
m = --------- = ------- = --- = 1
     x2 - x1     6 - 1     5

Then 

y - y1 = m(x - x1) becomes

 y - 2 = 1(x - 1)

 y - 2 = x - 1
   + 2     + 2
 -------------
 y     = x + 1

That's the equation:

     y = x + 1

Edwin